1. Let’s consider an AC circuit containing a coil. a) How does the impedance for this circuit change if the inductance is increased?
Answer #1

Impedance of an AC circuit is given by Z = \sqrt{}(R2 + (XL - XC)2)

=\sqrt{}(R2 + (wL - 1/wC)2)

Here we have considered only inductance L which means R = 0 and C = infinity

so Z = \sqrt{}(wL)2

= wL

we can see that Z is a function of inductance and frequency , so impedance increases with inductance.

So as the induction increases impedance also increases